Harold Johnson lives on a four-sided, 50,000-square-foot lot that is bounded on two sides by parallel streets. The city has assessed him $1,000 for curb repair, $2 for each foot of property bordering on these two streets. How far apart are the streets?
Answer by math_tutor2020(3835)


h = distance between parallel streets = height of trapezoid
b1, b2 = parallel sides

b1+b2 = total number of feet along curbside
2(b1+b2) = total cost = $1000
2(b1+b2) = 1000
b1+b2 = 1000/2
b1+b2 = 500

area of trapezoid = 0.5*h*(b1+b2) = 50,000
0.5*h*(b1+b2) = 50,000
0.5*h*(500) = 50,000
250h = 50,000
h = 50000/250
h = 200


Answer: 200 ft